• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Monaghan

    Log entry time 08:13:24 on March26,2005

    Entry number 142458

    keyword=Owl shift summary

    Owl Shift Summary

    Shift Workers
    Peter Monaghan
    Yi Jiang

    00:00 Shift begins with 40uA beam being delivered to the hall and run #2943 in progress.

    00:40 MCC called to ask if we had any restrictions on the beam recovery rate; there are none that I know of but I asked that MCC continue to deliver beam as is and not mess around with anything.

    01:50 Compton computer screen has hung up and I cannot access any of the compton controls; not sure how to proceed - emailed Sirish.

    02:45 Lost the compton stripchart which was running on adaql1; cannot ssh into the compton computer as the connection is refused.

    03:05 Upper neutron detector high voltage crate tripped off; unable to reset it remotely and so have asked MCC for a controlled access to power cycle the crate.

    03:25 Successfully power cycled high voltage crate in the hall and we are going back to beam permit now.

    03:45 Beam back on and we are taking data again.

    05:45 Hachemi happened to be reading the Halog (? - at this hour of the morning!!) and saw my entries about the compton computer. He came into the counting house, rebooted the compton machine and set up all the CODA and Epics guis again. We now have our compton stripcharts back again and the compton seems to be working well. Thanks to Hachemi!!

    06:30 Noticed that the electron and photon rates in the compton stripcharts had flatlined again. Tried stopping and restarting compton daq runs to see if it started the heartbeat again - seems to have worked for now.

    06:45 After restarting coda for the compton, the traces seem very spikey in nature.

    07:45 Talked to Sirish on the telephone concerning the spikey nature of the photon and electron traces and he explained that we are running the compton daq in 'production' mode, which is not quite ready for production yet - it does not do the integrate the signals correctly. Instead we should use the 'spyacq' configuration. I changed this in the compton daq, restarted another run and everything seems a little more stable now.

    08:00 Shift ends with run 2952 in progress.


    Run Summary

    40uA beam on the tilted carbon target.

    23:53 - 00:43 2943 200K events
    00:44 - 01:39 2944 200K events
    01:40 - 02:28 2945 200K events
    02:29 - 03:02 2946 134K events - ended early due to hv crate failure.
    03:44 - 04:31 2947 200K events
    04:32 - 05:18 2948 200K events
    05:19 - 06:04 2949 200k events
    06:04 - 06:52 2950 200K events
    06:52 - 07:41 2951 200K events
    07:42 2952 in progress